A metal oxide has formula m2o. It can thermally be decomposed to the metal and oxygen. 5.8g of the metal oxide forms 400mg of oxygen gas on complete decomposition. What is the atomic weight of the metal

Grade:11

1 Answers

Arun
6 years ago
The chemical equation for the thermal decomposition of the oxide is as follows:
  
          2M2O  ---------------->   4M    +        O2

Therefore 2 moles of the oxide produces 4 moles of the metal and 1 mole of oxygen in terms of ratio.

The actual mole of oxygen produced is:

moles = mass/molar mass
           =0.4/16
           = 0.025 moles
mole ratio between metal and oxygen = 4:1
therefore moles of the metal = 4 x 0.025 = 0.1 moles

The mass of the metal is 5.8 - 0.4 = 5.4 g
moles = mass/mm
mass= mm x moles
mm= mass/moles
   =  5.4/0.1
    =54